OpenStudy (anonymous):

Easy, I'm just lost :/ What is the solution to the equation log (3) x – log (3) 2 = 2 ? x = 18 x = 12 x = 9/2 x = 4/3

OpenStudy (anonymous):

You need to figure out how to represent 2 as a log with base 3. So Log(3)x=2 Well, 2^3 =8 so x is 8. You now have log (3) x - log (3) 2 = log (3) 8 You can take out the log(3) now. x-2=8 And I think you can get it from there.

OpenStudy (anonymous):

Thats not an answer choice though? I got x=10

OpenStudy (anonymous):

what? this makes no sense

OpenStudy (anonymous):

I'm an idiot. Log(3)x-Log(3)2=2 Log(3)(x/2)=2 3^2=(x/2) 9=(x/2) 18=x Don't even ask what I was thinking.

OpenStudy (anonymous):

Lol its fine :) thanks!